# Proactive Service Assurance in 5G and B5G Networks: A Closed-Loop Algorithm for End-to-End Network Slicing

### v3

@@ Abstract
The customization of services in Fifth-generation (5G) and Beyond 5G (B5G) networks relies heavily on network slicing, which creates multiple virtual networks on a shared physical infrastructure, tailored to meet specific requirements of distinct applications, using Software Defined Networking (SDN) and Network Function Virtualization (NFV). It is imperative to ensure that network services meet the performance and reliability requirements of various applications and users, thus, service assurance is one of the critical components in network slicing. One of the key functionalities of network slicing is the ability to scale Virtualized Network Functions (VNFs) in response to changing resource demand and to meet Customer Service Level agreements (SLAs).
In this paper, we introduce a proactive closed-loop algorithm for end-to-end network orchestration, designed to provide service assurance in 5G and B5G networks. We focus on dynamically scaling resources to meet key performance indicators (KPIs) specific to each network slice and operate in parallel across multiple slices, making it scalable and capable of managing completely automatically real-time service assurance. Through our experiments, we demonstrate that the proposed algorithm effectively fulfills service assurance requirements for different network slice types, thereby minimizing network resource utilization and reducing the over-provisioning of spare resources.
Index Terms:
5G/B5G resource allocation, 5G and B5G network slicing, closed-loop algorithm, end-to-end network slicing, proactive service assurance, quality of service, service assurance.
@@ I Introduction
5G and B5G cellular networks have introduced significant improvements in wireless communication technology, providing better transmission rates, ultra-low latency, and enhanced capacity. However, with the ever-increasing demand for low-latency and high-bandwidth applications, the current 5G architecture faces a range of challenges including resource allocation for network slicing, ensuring quality of service, and maintaining security and isolation among network slices, as reported in recent studies [1, 2]. These challenges must be effectively addressed in order to satisfy the changing requirements of both network operators and end-users. Furthermore, B5G networks will probably require networks with load adaptability and scalability to meet the growing demand for 5G and B5G applications. To meet the ever-increasing demand for high-speed, advanced reliability, and responsive data transmission, telecommunication networks have undergone significant transformations in recent years. Consequently, the wireless networks must support dynamic network slicing, VNF and SDN [3] to allow vendors to provide different kinds of services and fulfill the specific needs of each application type. The concept of network slicing, which enables a single physical network infrastructure to be partitioned into multiple virtual networks, has emerged as a promising solution to meet the distinct requirements of various applications and customers, see Zhang [4].
Likewise, to effectively manage and utilize network slices, the infrastructure should be able to dynamically allocate resources based on the service requirements of each slice. They include services with a wide variety of quality of service (QoS) needs, such as enhanced mobile broadband (eMBB), ultra-reliable low-latency communications (uRLLC), and massive machine-type communications (mMTC) [5]. In the context of VNF scaling within Service Assurance (SA), resource allocation is the process of allocating specific resources, such as CPU, memory, and storage, to VNF instances. This allocation can be done manually or automatically. Manual allocation requires more time and is more susceptible to errors, but it provides greater control over resource usage. Automated allocation can be more efficient, but it does not always allocate resources optimally. In particular, the VNF auto-scaling process involves balancing network actions and spare resources to meet the QoS, but also needs to achieve cost savings [6]. To guarantee QoS, a common practice is to wisely allocate additional resources to the network components to accommodate the traffic load demand. However, excessive allocation of resources can reduce cost savings. It is important to note that achieving the right balance between resource allocation and cost savings is crucial in the auto-scaling of VNF instances. Proper resource allocation can ensure that performance requirements are met, while cost savings can be achieved by avoiding excessive allocation of resources.
@@
To maintain optimal network performance and resource utilization, closed-loop control mechanisms must be implemented in network slicing [7]. The closed-loop algorithm is a feedback control system used in Service Assurance of wireless networks to improve network performance and maintain service quality. Closed-loop control mechanisms play a vital role in continuously monitoring performance and resource utilization, enabling real-time responses to satisfy the distinctive requirements of each slice in functional domains such as the radio access network (RAN), transport network (TN), and core network (CN). Achieving service assurance can involve modifying network configurations, better resource allocation, or improved management of traffic flows. For instance, if a slice needs more capacity, the control loop management and orchestration can assign additional resources to the slice in real-time without affecting the effectiveness of other slices. A closed-loop algorithm in the wireless network typically involves three steps:
  • Monitoring: Monitor the network performance and collect data on KPIs, such as network throughput, latency, and packet loss. This data is then analyzed to identify any issues that may be affecting network performance.
  • Analysis: Analyze the data and identify the root cause of any issues that are affecting the network performance. This may involve analyzing network traffic patterns, identifying bottlenecks, or diagnosing specific network components that are causing issues.
  • Control: Implement control mechanisms to address the issues identified in the analysis phase. This may involve adjusting network configurations, rerouting traffic, or implementing QoS measures to prioritize traffic.
@@
The utilization of a closed-loop algorithm for 5G and B5G network slicing yields numerous advantages. Implementing the mechanism can help guarantee optimal resources, network efficiency can be optimized. Additionally, network performance can be enhanced by reducing latency and jitter. Also, the network’s effectiveness can be improved by mitigating congestion and service disruptions. However, the development of a closed-loop algorithm also represents several challenges [8]. The network often has a high degree of complexity and is subject to various factors that may impact its operational efficiency. Furthermore, the addition of new hardware and applications causes the network to constantly evolve. The 5G network is made up of a variety of different components across the RAN, TN, and CN, which also might include cloud computing resources, making it heterogeneous. Thus, numerous research activities are currently in progress to devise closed-loop algorithms for network slicing in the 5G and B5G networks, organized by diverse entities such as academic institutions, industrial bodies, and governmental agencies. A closed-loop algorithm should possess certain characteristics such as:
@@
  • Security: The algorithm can protect the network against security risks, including denial-of-service attacks and slice isolation.
@@
This paper seeks to conduct an in-depth examination and enhancement of a scalable proactive closed-loop algorithm, named PCLANSA - Proactive Close Loop Algorithm for Network Slicing Assurance, with a focus on proactive characteristics for service assurance in 5G and B5G networks enabled with network slicing. The algorithm is designed to optimize (the best effort) the utilization of network resources while ensuring compliance with the QoS requirements of different slices. Moreover, PCLANSA is designed with flexible parameters which helps to adapt seamlessly to varying conditions and diverse network resources. This adaptability enhances its performance across various scenarios and contributes to its multi-functionality in addressing different QoS requirements. The paper is organized as follows. An overview of related work is presented in Section II, followed by a discussion of End-to-End (E2E) orchestration in Section III. We next present our E2E network slicing infrastructure. The design of our proposed proactive closed-loop algorithm for service assurance in 5G/B5G networks is detailed in Section IV. In Section V, we present an analysis of the efficiency of PCLANSA on E2E network slicing. Finally, we draw the conclusions in Section VI and discuss the benefits of this approach.
@@ II Related works
The management of service assurance in 5G and B5G networks is currently an important area of investigation. It implies the efficient management of network resources, and various algorithms have been proposed to dynamically allocate compute resources to VNF instances and manage network resources. The objective is to fulfill the SLAs while simultaneously minimizing resource utilization, operating costs, and energy consumption.
To effectively handle compute resources within 5G networks, Ren et al [9] proposed the use of NFV to virtualize network components in the core network, commonly referred to as the virtual Evolved Packet Core (vEPC). The authors also designed the Dynamic Auto Scaling Algorithm (DASA) for virtualizing the EPC in 5G networks. This algorithm uses a combination of queuing theory and cost analysis to balance the trade-off between performance and operation cost while considering the capacity of legacy network equipment. Then, several subsequent algorithms (see, e.g., [10]) have been proposed for dynamically scaling VNF instances in the vEPC in order to meet performance requirements while minimizing resource usage. Ren et al [11] also proposed the Adaptive VNF Scaling Algorithm (ASA), designed to balance the cost-performance trade-off in 5G mobile networks. This algorithm uses an analytical model to determine the optimal number of VNF instances needed to handle data traffic while minimizing the cost.
While the aforementioned algorithms demonstrate considerable potential in addressing the challenges of dynamically scaling VNF instances, their primary focus remains within one network domain, with limited configurable parameters. Furthermore, their efficiency is heavily dependent on the current traffic load, without taking into account other factors such as delay, jitter, packet loss, etc. This makes them less adaptable to the varying requirements posed by diverse network applications. For instance, accommodating the distinct requirements of applications such as eMBB and mMTC might prove challenging.
Considering the literature on cellular networks, various categories of ”slicing problems” have received attention and exploration. One of the central areas that emerge is the allocation challenge resources for physical nodes between slices, including allocation of resource blocks within the CN and RAN [12, 13, 14, 15, 16]. In terms of 5G network slicing, VNF refers to software-based network functions (instantiated as virtual machines or containers) that operate on virtualized infrastructure, delivering services and functionality within network slices. The challenge of VNF resource allocation in SA mainly focuses on optimizing the utilization of existing resources while satisfying the various requirements of distinctive network slices. The requirements include but are not limited to:
@@
  • QoS requirements: network slices may exhibit diverse QoS requirements, that include factors such as latency, throughput, reliability, and availability. In order to fulfill the service-level agreements (SLAs) for each slice, it is essential that resource allocation takes into account these particular requirements.
@@
  • Dynamic resource demands: the demands for resources may vary dynamically depending on factors such as network traffic patterns, user behaviours, and application requirements. Thus, the adaptability and real-time responsiveness of the resource allocation mechanism is important.
@@
  • Multi-dimensional resource optimization: the process of resource allocation within the context of 5G network slicing entails the simultaneous optimization of various dimensions, including but not restricted to CPU utilization, memory usage, power consumption, and network bandwidth. Therefore, the optimization problem for satisfying the QoS requirements for multiple slices while maintaining a balance between the different dimensions is a complex challenge.
@@
  • Isolation and security: ensuring proper isolation between slices is crucial to prevent interference, unauthorized access, and data breaches. Additionally, maintaining robust security measures within each slice is imperative to protect sensitive information and mitigate potential vulnerabilities.
@@
  • Resource allocation policies: to ensure optimal performance, it is essential to design resource allocation policies that effectively allocate resources based on the specific needs (such as those derived from SLA) and priorities of each slice. This requires considering factors such as QoS requirements, traffic demands, latency constraints, and dynamic resource allocation.
Some research efforts use closed-loop mechanisms to address network slicing AS challenges [17, 18, 19]. The majority of the current research on slice embedding mainly concentrates on addressing the one-shot optimization problem, which involves optimizing resource allocation based on average and/or static demands. The primary goal of SA is to dynamically and real-time allocate resources to VNF instances or network installations in order to meet SA requirements while minimizing resource usage.
In summary of the above analyses, the aforementioned articles highlight some limitations, notably the lack of consideration of network KPIs based on the various QoS requirements of different slices. Our study aims to fulfill these limitations and therefore determine an efficient approach to resource allocation while using a forecasting model. In particular, it will take into account a wider range of factors, including the data rate of the link, the concept of E2E network slicing and the distinct quality of service of each of the slices.
@@ III End-to-End Orchestration
With the new design mentioned above, the concept of E2E orchestration has recently gained prominence as an innovative concept in the domain of 5G and B5G networks [21]. Orchestration refers to the comprehensive management as well as coordination of multiple network functions, resources and services across the network infrastructure, resulting in parallel degrees of flexibility, efficiency and operation automated. Demonstrated in Fig 1 from 3GPP, the implementation of end-to-end orchestration provides a holistic strategy for handling network operations, supporting operators to efficiently manage and enhance all network components, ranging from the RAN, the TN to the CN. The E2E network slicing infrastructure employs various management domains and utilizes modern SDN and NFV technologies to facilitate flexible resource allocation, service chaining, and policy enforcement. Thus, E2E network slicing involves a physical infrastructure comprising network, computing, and storage resources that are programmable and embedded throughout the end-to-end communication paths.
Figure 1: A reference E2E network slicing concept [22]
The studies [23, 24] provide a deep overview of E2E network slicing in both vertical and horizontal directions with a detailed discussion on slice isolation, and application use cases that enable a comprehensive infrastructure for network slicing in a 5G network. Based on the showcases, we can indicate the significance of slice isolation, which guarantees the independent and secure operation of each network slice, without any external factors or interference from other slices. Ensuring the confidentiality, integrity, and efficiency of each slice is of the highest priority in situations where sensitive or vital applications are utilized, thereby emphasizing the significance of slice isolation. Thus, the mechanisms and techniques need to be revisited to create slice isolation effectively and to address various challenges that arise in this scenario, including resource allocation, traffic management, and security enforcement. As per the definition provided in reference [25], the concept of network slicing is comprised of three distinct layers.
@@
Despite the numerous benefits that E2E network slicing offers for 5G and B5G networks, there remain certain gaps in knowledge and research opportunities [26] such as RAN virtualization and slicing, holistic and intelligent slice orchestration, secure sliced networks, quality of services in network slicing. Drawing from the previously mentioned review, below we will construct a 5G E2E network slicing architecture in a simulator environment, with the goal of implementing and addressing intelligent network management in the context of service assurance.
@@
Fig. 2 depicts a high-level view of our E2E network slicing infrastructure with a closed-loop algorithm on a 5G network.
It consists of five primary components, including slice control, MANagement and Orchestration (MANO), virtualized networks/platforms, physical infrastructure, and a closed-loop algorithm. In the virtualized networks/platform, there exists a set of commonly shared network functions (NFs) [27, 28], including but not limited to the Network Slice Selection Function (NSSF), Policy Control Function (PCF), and Access and Mobility Management Function (AMF). This approach offers several advantages, such as cost savings on hardware and software, enhanced network efficiency via a reduction in the number of VNF instances that must be deployed, and increased network scalability by facilitating the creation of new network slices. In addition, slice control is used to establish and manage network slices, enforce slice policies, and monitor slice performance. In cooperation with slice control, the MANO component is in charge of ensuring optimal network performance and functionality. This includes facilitating network visibility, equipping network administrators with effective management tools, and automating network management processes [29]. Next, the integration of a closed-loop algorithm has been implemented with the goal of improving coordination between the slice control and MANO components. This integration has enabled the management of network slicing in reliable and efficient ways, while also aligning with customer requirements and satisfying QoS. Finally, the aforementioned components are in charge of controlling and managing a shared physical infrastructure in order to establish an E2E network that optimizes the entirety of the network capability, from the RAN, the TN and the CN. Thus, every E2E network slice is created with an isolated virtual network, a set of VNF instances, dedicated virtual computing and storage resources with several shared common NFs.

### v4

# Proactive Service Assurance in 5G and B5G Networks: A Closed-Loop Algorithm for End-to-End Network Slices

@@ Abstract
Ensuring the highest levels of performance and reliability for customized services in fifth-generation (5G) and beyond (B5G) networks requires the automation of resource management within network slices. In this paper, we propose PCLANSA, a proactive closed-loop algorithm that dynamically allocates and scales resources to meet the demands of diverse applications in real time for an end-to-end (E2E) network slice. In our experiment, PCLANSA was evaluated to ensure that each virtualized network function is allocated the precise resources it requires, thereby maximizing efficiency and minimizing waste. This goal is achieved through the intelligent scaling of virtualized network functions. The benefits of PCLANSA have been demonstrated across various network slice types, including eMBB, mMTC, uRLLC, and VoIP. This finding indicates the potential for substantial gains in resource utilization and cost savings, with the possibility of reducing over-provisioning by up to 54.85%.
@@ I Introduction
The increasing demand for diverse, high-performance applications in 5G and B5G networks necessitates efficient resource allocation and service assurance within network slices. While network slicing offers customized service delivery, ensuring that each network slice meets its performance requirements (e.g., low latency for uRLLC, high throughput for eMBB) while minimizing resource consumption presents a significant challenge. Existing closed-loop service assurance mechanisms often react to performance degradations, leading to potential Service Level Agreements (SLAs) violations and inefficient resource utilization, as highlighted in recent studies [1, 2]. To meet the evolving requirements of both network operators and end-users, these challenges must be effectively addressed. In particular, B5G networks will demand greater load adaptability and scalability to support the rapid growth of 5G and B5G applications. As a result, telecommunication networks have undergone substantial transformations in recent years to deliver higher speeds, enhanced reliability, and more responsive data transmission. Within this context, machine learning plays a pivotal role by enabling proactive network behaviour through real-time prediction, anomaly detection, and intelligent decision-making.
In network slicing, the infrastructure must demonstrate the capacity to dynamically allocate resources in accordance with the service requirements of each network slice. These services encompass a broad spectrum of quality of service (QoS) needs, as detailed in the reference [3]. In order to meet the QoS requirements, resource allocation is the common process of allocating specific resources, such as central processing units (CPUs), memory, and storage, to virtual network functions (VNF) instances. This allocation can be executed manually or automatically. Manual allocation requires a greater investment of time and is more susceptible to errors, but it provides better control over resource usage. Automated allocation can be more efficient but may not always allocate resources optimally. Thus, the VNF auto-scaling process entails a delicate balancing act between network actions and spare resources, with the objective of meeting QoS requirements while achieving cost savings, as articulated in the work of Rahman et al. [4].
@@
Network performance and resource utilization are often optimized with closed-loop control mechanisms in network slice [5]. The closed-loop algorithm is a feedback control system used in Service Assurance (SA) of communication networks to improve network performance and maintain service quality. Closed-loop control mechanisms play a vital role in continuously monitoring performance and resource utilization, enabling real-time responses to satisfy the distinctive requirements of each network slice in functional domains such as the radio access network (RAN), transport network (TN), and core network (CN). Achieving service assurance can involve modifying network configurations, better resource allocation, or improved management of traffic flows. For instance, if a network slice needs more capacity, the control loop management and orchestration system can assign additional resources to the network slice in real-time without affecting the effectiveness of other network slices.
@@
The employment of a closed-loop algorithm for 5G and B5G network slices yields numerous advantages, including the optimization of resources and network efficiency, as well as enhanced network performance through reduced latency and jitter. Furthermore, network efficiency can be increased by mitigating congestion and service disruptions. However, the development of a closed-loop algorithm also poses several challenges, as outlined in [6]. The network often has a high degree of complexity and is subject to various factors that can affect its operational efficiency. Thus, numerous research activities are currently in progress to devise closed-loop algorithms for network slices in the 5G and B5G networks, organized by various entities such as academic institutions, industrial bodies, and government agencies. A closed-loop algorithm should possess certain characteristics, such as:
@@
  • Security: The algorithm can protect the network against security risks, including denial-of-service attacks and network slice isolation.
@@
The present study aims to propose an in-depth examination and enhancement of a scalable proactive closed-loop algorithm, PCLANSA - Proactive Close Loop Algorithm for Network slice Assurance, with a focus on proactive characteristics for service assurance in 5G and B5G networks enabled with network slices. The algorithm is designed to optimize the utilization of network resources while ensuring compliance with the QoS requirements of multiple network slices operating in parallel. Additionally, the PCLANSA is designed with flexible parameters that facilitate seamless adaptation to variable conditions and diverse network resources, thereby enhancing its performance across a range of scenarios and contributing to its multi-functionality in addressing disparate QoS requirements.
The remainder of this paper is organized as follows. Section II reviews related work on service assurance and resource allocation in network slices. Section III provides a concise overview of the E2E network slice architecture. Section IV details the design and implementation of PCLANSA. Section V presents a comprehensive performance evaluation of PCLANSA using a realistic simulation environment. Finally, Section VI concludes the paper, discusses the benefits and future research directions.
@@ II Related works
The automation of resource management and service assurance in next-generation networks has been extensively studied, particularly in the context of 5G and network slicing. This entails the efficient management of network resources, and numerous algorithms have been proposed to dynamically allocate compute resources to VNF instances while optimizing network performance. The primary objective is to meet SLAs while simultaneously minimizing resource utilization, operational costs, and energy consumption. A central challenge lies in the dynamic and intelligent allocation of resources to ensure QoS across heterogeneous and customized network slices.
Early approaches to service assurance primarily relied on manual configuration and reactive mechanisms. However, the increasing complexity of modern networks has necessitated the adoption of closed-loop automation, in which systems can autonomously monitor, analyze, plan, and execute (MAPE) actions. To efficiently manage compute resources in 5G networks, several algorithms and frameworks have been developed to extend the MAPE loop. For example, Ren et al. [7] proposed a distributed closed-loop architecture for real-time orchestration and service assurance. This work emphasizes a hierarchical control plane and a knowledge-based service assurance system. In addition, the DASA algorithm in their work addresses dynamic resource allocation for 5G network slices. Similarly, Ali et al. [8] introduced a service-assurance-based closed-loop framework for managing virtualized networks. These methods, however, often require that a Key Performance Indicator (KPI) threshold be violated before any corrective action is taken, which is insufficient for applications with stringent latency and reliability requirements. Other notable works, such as the Adaptive Service Assurer (ASA) [9] and Govindarajan et al. [10], also rely on reactive adjustments to maintain service levels. While these frameworks optimize resource allocation, they predominantly operate in a reactive mode, responding only after a performance issue has occurred. This reactive nature limits their capacity to anticipate fluctuations in network demand or proactively prevent service degradation. Therefore, there is a clear need for predictive and intelligent mechanisms that can dynamically allocate resources in advance, ensuring continuous compliance with QoS requirements and supporting the strict performance demands of modern 5G and beyond networks.
Considering the literature on cellular networks, various categories of ”slicing problems” have received attention and exploration. One of the central areas that emerges is the allocation challenge resources for physical nodes among network slices, including allocation of resource blocks within the CN and RAN [11, 12]. The complexity of VNF resource allocation in SA is primarily characterized by the optimization of existing resources in a manner that satisfies the diverse requirements of distinct network slices [13]. These requirements encompass, but are not limited to, the following:
@@
  • QoS requirements: network slices may exhibit diverse QoS requirements, which include factors such as latency, throughput, reliability, and availability. In order to fulfill the SLAs for each slice, it is essential that resource allocation takes into account these particular requirements.
@@
  • Dynamic resource demands: the demands for resources may vary dynamically depending on factors such as network traffic patterns, user behaviours, and application requirements. Thus, the adaptability and real-time responsiveness of the resource allocation mechanism are important.
@@
  • Multi-dimensional resource optimization: the process of resource allocation within the context of 5G network slice entails the simultaneous optimization of various dimensions, including but not restricted to CPU utilization, memory usage, power consumption, and network bandwidth. Therefore, the optimization problem for satisfying the QoS requirements for multiple slices while maintaining a balance between the different dimensions is a complex challenge.
@@
  • Isolation and security: ensuring proper isolation between network slices is crucial to prevent interference, unauthorized access, and data breaches. Additionally, maintaining robust security measures within each network slice is imperative to protect sensitive information and mitigate potential vulnerabilities.
@@
  • Resource allocation policies: to ensure optimal performance, it is essential to design resource allocation policies that effectively allocate resources based on the specific needs (such as those derived from SLA) and priorities of each network slice. This requires considering factors such as QoS requirements, traffic demands, latency constraints, and dynamic resource allocation.
In the context of network slicing, certain research endeavours have employed closed-loop mechanisms to address the associated challenges, such as [14, 15]. The majority of existing research on network slice embedding has focused on addressing the one-shot optimization problem, which involves optimizing resource allocation based on average and/or static demands. However, the latest evolution of the primary objective of SA is to dynamically and in real time allocate resources to across network slices or network installations in order to meet SA requirements while minimizing resource usage. Furthermore, a critical limitation in much of the existing literature is a focus on isolated network segments. Many proposals address resource management within the Radio Access Network (RAN) or core network (CN) or the transport network (TN) in isolation [16, 17, 18]. This fragmented approach fails to account for the holistic, end-to-end (E2E) performance of a network slice, where performance bottlenecks can arise at any point along the service chain. An effective solution must be capable of orchestrating resources across the entire E2E path to guarantee a seamless and consistent user experience.
In contrast to these existing efforts, our proposed PCLANSA introduces a novel approach that overcomes these limitations. While prior closed-loop systems are predominantly reactive, PCLANSA is inherently proactive. It leverages a forecasting model to anticipate future resource demands and potential network congestion, enabling the system to scale resources before performance degradation occurs. This predictive capability allows PCLANSA to maintain high levels of QoS, even under rapidly changing network loads. Moreover, unlike solutions focused on individual network segments, PCLANSA provides end-to-end network orchestration. Our work advances the field of network slice assurance by developing a proactive closed-loop algorithm that integrates machine learning for dynamic resource allocation across end-to-end 5G network slices. In contrast to Marinova et al. [19], whose research presents a holistic framework for E2E network slice assurance through data collection, MLOps, and multi-domain closed-loop control with a focus on system architecture and operational workflows, our approach emphasizes algorithmic innovation for predictive scaling and SLA adherence in dynamic network environments. This focus enables real-time adaptation to traffic variations within network slices, reducing KPI violations and optimizing resource utilization. It intelligently allocates and scales resources across the entire E2E network slice, including both the core and transport domains. By adopting this holistic perspective, PCLANSA ensures service assurance across the entire service chain, making it more robust and effective in managing the complexities of modern, virtualized 5G and B5G networks. The comparative analysis presented in our evaluation section will further highlight the significant performance improvements achieved by our proactive and end-to-end approach.
@@ III End-to-End Orchestration
With the new design mentioned above, the concept of E2E orchestration has recently gained prominence as an innovative concept in the domain of 5G and B5G networks [21]. Orchestration refers to the comprehensive management as well as coordination of multiple network functions, resources and services across the network infrastructure, resulting in parallel degrees of flexibility, efficiency and automated operation. Demonstrated in Fig. 1 from 3GPP, the implementation of end-to-end orchestration provides a holistic strategy for handling network operations, supporting operators to efficiently manage and enhance all network components, ranging from the RAN, the TN, to the CN. The E2E network slice infrastructure employs various management domains and utilizes modern SDN and NFV technologies to facilitate flexible resource allocation, service chaining, and policy enforcement. Thus, an E2E network slice involves a physical infrastructure comprising network, computing, and storage resources that are programmable and embedded throughout the end-to-end communication paths.
Figure 1: A reference E2E network slice concept [22]
The study [23] provides a deep overview of E2E network slice in both vertical and horizontal directions with a detailed discussion on network slice isolation, and application use cases that enable a comprehensive infrastructure for network slice in a 5G network. Based on the showcases, we can indicate the significance of network slice isolation, which guarantees the independent and secure operation of each network slice, without any external factors or interference from other network slices. Ensuring the confidentiality, integrity, and efficiency of each network slice is of the highest priority in situations where sensitive or vital applications are utilized, thereby emphasizing the significance of network slice isolation. Thus, the mechanisms and techniques need to be revisited to create network slice isolation effectively and to address various challenges that arise in this scenario, including resource allocation, traffic management, and security enforcement. As per the definition provided in reference [24], the concept of network slice comprises three distinct layers.
@@
Despite the numerous benefits that E2E network slice offers for 5G and B5G networks, there remain certain gaps in knowledge and research opportunities [25] such as RAN virtualization and network slice, holistic and intelligent network slice orchestration, secure network slices, and quality of services in multiple network slices. Drawing from the previously mentioned review, below we will construct a 5G E2E network slice architecture in a simulator environment, with the goal of implementing and addressing intelligent network management in the context of service assurance.
@@
Fig. 2 depicts a high-level view of our E2E network slice infrastructure with a closed-loop algorithm on a 5G network.
It consists of five primary components, including network slice control, MANagement and Orchestration (MANO), virtualized networks/platforms, physical infrastructure, and a closed-loop algorithm. In the virtualized networks/platform, there exists a set of commonly shared network functions (NFs) [26, 27], including but not limited to the Network Slice Selection Function (NSSF), Policy Control Function (PCF), and Access and Mobility Management Function (AMF). This approach offers several advantages, such as cost savings on hardware and software, enhanced network efficiency via a reduction in the number of VNF instances that must be deployed, and increased network scalability by facilitating the creation of new network slices. In addition, network slice control is used to establish and manage network slices, enforce network slice policies, and monitor slice performance. In cooperation with network slice control, the MANO component is in charge of ensuring optimal network performance and functionality. This includes facilitating network visibility, equipping network administrators with effective management tools, and automating network management processes [28]. Next, the integration of a closed-loop algorithm has been implemented with the goal of improving coordination between the network slice control and MANO components. This integration has enabled the management of network slice in reliable and efficient ways, while also aligning with customer requirements and satisfying QoS. Finally, the aforementioned components are in charge of controlling and managing a shared physical infrastructure in order to establish an E2E network that optimizes the entirety of the network capability, from the RAN, the TN and the CN. Thus, every E2E network slice is created with an isolated virtual network, a set of VNF instances, dedicated virtual computing and storage resources, with several shared common NFs.